# Gradient entropy (GradEn): The two dimensional version of slope entropy for image analysis

### v1

@@ Abstract
Information theory and Shannon entropy are essential for quantifying irregularity in complex systems or signals. Recently, two-dimensional entropy methods, such as two-dimensional sample entropy, distribution entropy, and permutation entropy, have been proposed for analyzing 2D texture or image data. This paper introduces Gradient entropy (GradEn), an extension of slope entropy to 2D, which considers both symbolic patterns and amplitude information, enabling better feature extraction from image data. We evaluate GradEn with simulated data, including 2D colored noise, 2D mixed processes, and the logistic map. Results show the ability of GradEn to distinguish images with various characteristics while maintaining low computational cost. Real-world datasets, consist of texture, fault gear, and railway corrugation signals, demonstrate the superior performance of GradEn in classification tasks compared to other 2D entropy methods. In conclusion, GradEn is an effective tool for image characterization, offering a novel approach for image processing and recognition.
Index Terms:
Gradient entropy, Image processing, Texture analysis, Fault diagnosis, Irregularity
I Introduction
@@
Entropy-based algorithms have emerged as fundamental tools for quantifying the irregularity and uncertainty inherent in time series analysis and complex signal characterization[1, 2, 3, 4]. Originating from Shannon’s information theory[5], these methods operate by analyzing the probability distribution of distinct states within temporal data, thereby revealing hidden structural patterns in dynamical systems. Over recent decades, researchers have progressively developed various entropy quantification techniques to address different analytical challenges. Notable advancements include approximate entropy[6], sample entropy[7], fuzzy entropy[8], distribution entropy[9], permutation entropy[10], and dispersion entropy[9], each offering unique perspectives on system complexity assessment. The practical effectiveness of these entropy measures has been demonstrated across diverse application domains, including biological signal processing, financial market analysis, and industrial system monitoring, where they exhibit remarkable capabilities in differentiating complex signal patterns.
@@
Recent years have witnessed a significant expansion of information theory and entropy-based methodologies into texture and image analysis. Building upon their success in one-dimensional signal processing, two-dimensional entropy measures have emerged as powerful tools for quantifying texture complexity by capturing the spatial irregularity and uncertainty inherent in image data. This methodological evolution has led to the development of specialized 2D entropy algorithms, including two dimensional sample entropy ($SampEn_{2d}$) [11], two dimensional distrbution entropy ($DistrEn_{2d}$) [12], two dimensional permutation entropy ($PerEn_{2d}$) [13] and two dimensional dispersion entropy ($DispEn_{2d}$) [14]. all of which have demonstrated efficacy in analyzing both synthetic textures and real-world images.
@@
In this study, we introduce Gradient Entropy (GradEn), a novel two-dimensional extension of Slope Entropy (SloE)[15], specifically designed for analyzing 2D datasets. Building on the strengths of its one-dimensional counterpart, GradEn uniquely combines both symbolic representations and amplitude variations at each spatial point. This integration enhances GradEn’s discriminative power, allowing for more effective characterization of diverse image data and its classification applications. We begin by discussing the selection of parameters and the robustness of the GradEn method when applied to 2D white noise. Next, we conduct experiments on simulated 2D data, including colored noise and mixed processes, to validate its classification capability. For logistic mapping, we employ the distance matrix from the recurrence plot method to convert the 1D time series into an image, enabling the identification of distinct images generated by different logistic parameters. Additionally, we discuss the computational cost, demonstrating the efficiency of GradEn. The method is also applied to real-world data: two real texture datasets are selected for classification tasks, and GradEn is used to detect and diagnose gear faults and railway corrugation in real-world signals after transforming these signals into matrix form. In both simulated and real-world experiments, the results of the proposed GradEn method are compared with other state-of-the-art 2D entropy algorithms, including $SampEn_{2d}$, $DispEn_{2d}$ and $PerEn_{2d}$. The classification results highlight the superiority and effectiveness of GradEn in distinguishing different types of textures and images, providing a novel perspective for analyzing the properties and structures of 2D data.
@@
The content of this paper is listed as follows: Section II provides a detailed explanation of the procedure for calculating GradEn. Section III presents the results of applying GradEn to simulated 2D data, demonstrating the effectiveness of the proposed method. In Section IV, we use real-world texture datasets and signals for classification using GradEn. Finally, Section V concludes the paper.
II Gradient Entropy
@@
Slope entropy, introduced by [15] in 2015, is based on calculating the slope between consecutive points in a time series. This method offers a unique approach by incorporating both the symbolic pattern and the amplitude information within the series. When applied to 2D images, the concept of slope can be extended and replaced by the gradient of pixels in various directions.
@@
Assume that we have an image $X=\{x_{i,j}\}_{i=1,2,...,H}^{j=1,2,...,W}$, where the height equals to $H$ and width equals to $W$. The sketch map of GradEn algorithm is shown in Fig. 1 and the procedure is listed as follows:
@@
III Experiments in simuated data
III-A Parameter Analysis
First of all, the selection of pixel block dimensions ($m*n$) in GradEn requires systematic analysis, similar to the determination of embedding dimension $m$ in SloE. While the block size could theoretically be arbitrary, practical implementation must consider the structural characteristics of natural images. Smaller configurations such as $1*2$ or $2*1$ blocks yield only $55^{1}=$ 5 distinct gradient patterns(see Fig. 2(b)), which proves insufficient for effective information characterization. Conversely, larger configurations like $2*3$ or $3*2$ blocks generate $6255^{6}=15,$ 625 potential patterns (see Fig. 2(c)), creating dual challenges of computational complexity and noise sensitivity. Through evaluation, we identify the $2\texttimes 2$ configuration as an optimal compromise: it produces $1255^{3}=$ 125 distinctive patterns - a quantity that maintains discriminative power while ensuring computational efficiency and noise robustness.
@@ III-A Parameter Analysis
Then, the thresholds $\delta$ and $\gamma$ should be selected carefully. Due to the z-score standardization, the range of $\delta$ is constrainted to $(\mu_{0.5},\mu_{0.75})$ and $\gamma$ falls within $(\mu_{0.75},\mu_{1})$, where $\mu_{\alpha}$ represents the $\alpha$ quantile of the normal distribution. Let $\delta=\mu_{a}$ and $\gamma=\mu_{b}$, a 100*100 2D whitenoise image with the size of 100*100 is then generated, and the GradEn values for $a$ changing from 0.51 to 0.74 and $b$ changing from 0.76 to 0.95, with step equals to 0.01, are shown in Fig. 3. Given the high degree of irregularity in the white noise data, the entropy value is expected to be relatively large. Therefore, the parameter values in the yellow region ($a\in$ [0.53,0.6], $b\in$ [0.8,0.86]) in Fig. 3 are recommended. In this paper, $a=0.55$ and $b=0.8$ as the thresholds.
@@
III-B Colored noise classification
@@ III-C Robust Analysis
III-D Logistic map
The logistic map is a widely used chaotic model due to its ability to exhibit a variety of nonlinear behaviors as the parameter changes[16]. The formulation of the logistic map is given by:
@@ III-D Logistic map
where $a$ is the control parameter that determines the degree of periodicity or chaos in the sequence. To assess the ability of GradEn to distinguish between different states in the logistic map, we transform the logistic sequence into a distance matrix, similar to the preprocessing algorithm used in recurrence plots. Given the sequence $X_{i},i=1,2,...,N$, first perform phase space reconstruction with an embedding dimension $m$, resulting in a series of vectors $x(i)=(X_{i},X_{i+1},...,X_{i+m-1})$ for $i=1,2,...,N-m+1$. Next, construct the distance matrix $D=D_{i,j}$, where $i=1,2,...,N-m+1$ and $j=1,2,...,N-m+1$, with each element defined as $D_{i,j}=dist(x(i),x(j))$. The Euclidean distance is used in this algorithm. Once the matrix $D$ is constructed, it is then processed using the 2D entropy method to calculate the results
@@ III-E Computation cost
TABLE I: The Hedges’ g effect size results for the four methods across 15 EEG channels
@@
IV Experiments in real-world data
IV-A Texture Classification
Texture analysis is a key research area in image processing, focused on extracting the characteristics of various patterns within images. In this section, two texture datasets are used to evaluate the classification performance of the proposed $GradEn$ method. We begin by conducting experiments on the Kylberg texture dataset[17], which includes five selected texture categories: cushion, floor, scarf, screen, and stone. Each category contains 160 samples, and examples of the images are shown in Fig. 9. To standardize image size and ensure computational efficiency, all images are downsampled to 128×128. The boxplot results for $DistrEn_{2d}$ ($M=64$ and $M=128$), $PerEn_{2d}$, and $GradEn$ are presented in Fig. 10. As shown in Fig. 10(a) and (b), $DistrEn_{2d}$ fails to effectively distinguish between the five texture types, particularly for cushion, floor, and scarf. Similarly, $PerEn_{2d}$ ([Fig. 10(c)]) struggles to differentiate between cushion vs. scarf and floor vs. screen. In contrast, $GradEn$ demonstrates a clear ability to differentiate all five textures, as shown in [Fig. 10(d)], highlighting the superiority of the $GradEn$ method.
@@ IV-A Texture Classification
Additionally, the Describable Texture Dataset (DTD)[18] is used for the classification task. Four texture types—knitted, polka, wrinkled, and zigzagged—are selected, with 120 samples for each type and the examples are shown in Fig. 11. All images are grayscaled processed by taking the means of RGB values and also downsampled to 128×128 for consistency. The boxplot results for the four entropy algorithms are presented in Fig. 12. It is evident that only $GradEn$ is able to distinguish between the four texture groups to some extent, while the other three methods fail to differentiate the texture types effectively. In conclusion, both two experiments showcase the effectiveness of $GradEn$ for detecting the characteristic of texture images.
@@
IV-B Gear fault diagnosis
Mechanical fault diagnosis plays a crucial role in industrial engineering, and numerous methods have been proposed by researchers in recent years [19, 20, 21, 22]. In this section, gear fault data from the University of Connecticut are used for classification. The vibration signals are collected from a benchmark two-stage gearbox with replaceable gears. These signals are measured using an accelerometer and recorded through a dSPACE system, with a sampling frequency of 20 kHz. Four fault conditions—healthy, missing tooth, spalling, and chipping rip—are selected for the classification task (The examples are shown in Fig. 13). Each fault type consists of 104 samples. The first 150 points of each sample are used to construct the distance matrix with an embedding dimension of $m=3$. The boxplot results for classifying the four gear faults are shown in Fig. 14. As seen in the figure, $DistrEn_{2d}$ fails to effectively distinguish between the different fault types. $PerEn_{2d}$ struggles to differentiate between healthy vs. missing tooth and spalling vs. chipping rip. In contrast, only the $GradEn$ method is able to distinguish all four gear fault types based on entropy values, demonstrating its capability to characterize images generated by engineering signals.
@@ IV-B Gear fault diagnosis
IV-C Railway Corrugation Diagnosis
The detection and diagnosis of railway corrugation is one of the most important technology part for keeping the security of train operation[23, 24, 25]. In this section, the acceleration signals collected from the sections where corrugation exists of heavy haul railway are applied to verify the detection ability of $GradEn$. Fig. 15(a) and Fig. 15(b) gives the two acceleration signals we used which collected from the left and right axes of the railway section. The corrugation sections are marked by red frame in the picture. In our experiment, the signals with the length of $N=4000$ in both corrugation section (red frame) and normal sections (blue frame) are selected. The sliding window with $N=150$ with the step equals to 10 moves from the beginning to the end of the section. Thus, totally 385 samples of sections from both corrugation and normal signals are obtained. Then, all signals are transformed to distance matrix with embedding dimension $m=3$. The boxplot results of three entropy methods ($DistrEn_{2d}$, $PerEn_{2d}$, $GradEn$) for distinguishing corrugation and normal signals in both left and right axes are shown in Fig. 15(c-e) and (f-h). Moreover, the coresponding hedge effect size are calculated and given in each picture. The figures and numerical values manifest that $GradEn$ holds the best distinguishing results in both left and right axes signals, which proves that $GradEn$ method provides an effective tool for railway diagnosis research.
@@ V Conclusion
CRediT authorship contribution statement
Runze jiang: Conceptualization; Data Curation; Formal analysis; Methodology; Software; Validation; Visualization; Writing/Original draft preparation; Writing/Review & Editing.
Pengjian Shang: Formal analysis; Funding Acquisition; Methodology; Visualization; Writing/Review & Editing.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.
Acknowledgment
The financial supports from the Fundamental Research Funds for the Central Universities, China (2024YJS164) and the National Natural Science Foundation of China (62171018).

### v2

# Gradient Entropy (GradEn): The Two-dimensional Version of Slope Entropy for Image Analysis

@@ Abstract
Information theory and Shannon entropy are essential for quantifying irregularity in complex systems or signals. Recently, two-dimensional entropy methods, such as two-dimensional sample entropy, distribution entropy, and permutation entropy, have been proposed for analyzing 2D texture or image data. This paper introduces Gradient entropy (GradEn), an extension of slope entropy to 2D, which considers both symbolic patterns and amplitude information, enabling better feature extraction from image data. We evaluate GradEn with simulated data, including 2D colored noise, 2D mixed processes, and the logistic map. Results show the ability of GradEn to distinguish images with various characteristics while maintaining low computational cost. In classification tasks involving real-world datasets - including texture analysis, fault gear diagnosis, and railway corrugation detection - GradEn also exhibits significantly better performance than conventional 2D entropy methods. In conclusion, GradEn is an effective tool for image characterization, offering a novel approach for image processing and recognition.
keywords:
[…]
1 Introduction
@@
Entropy-based algorithms have emerged as fundamental tools for quantifying the irregularity and uncertainty inherent in time series analysis and complex signal characterization (Huang et al. (2012); He et al. (2021); Yang et al. (2023); Jiang et al. (2025)). Originating from Shannon’s information theoryShannon (1948), these methods operate by analyzing the probability distribution of distinct states within temporal data, thereby revealing hidden structural patterns in dynamical systems. Over recent decades, the progressive development of entropy quantification techniques has addressed diverse challenges in real world applications, with prominent methodologies including: approximate entropyYan and Gao (2007), sample entropyRichman and Moorman (2000), fuzzy entropy (Chen et al. (2009)), permutation entropy (Christoph and Bernd (2002)), distribution entropy (Li et al. (2015)) and dispersion entropy (Rostaghi and Azami (2016)). Each methodology provides distinct mathematical frameworks for quantifying system complexity. The practical effectiveness of these entropy measures has been demonstrated across diverse application domains, including biological signal processing, financial market analysis, and industrial system monitoring, where they exhibit remarkable capabilities in differentiating complex signal patterns.
@@
Recent years have witnessed a significant expansion of information theory and entropy-based methodologies into texture and image analysis. Building upon their success in one-dimensional signal processing, two-dimensional entropy measures have emerged as powerful tools for quantifying texture complexity by capturing the spatial irregularity and uncertainty inherent in image data. This methodological evolution has led to the development of specialized 2D entropy algorithms, including two dimensional sample entropy ($SampEn_{2d}$) (Silva et al. (2016)), two dimensional distrbution entropy ($DistrEn_{2d}$) (Azami et al. (2017)), two dimensional permutation entropy ($PerEn_{2d}$) (Gaudencio et al. (2022)) and two dimensional dispersion entropy ($DispEn_{2d}$) (Azami (2019)). all of which have demonstrated efficacy in analyzing both synthetic textures and real-world images.
@@
In this study, we introduce Gradient Entropy (GradEn), a novel two-dimensional extension of Slope Entropy (SloE) (Cuesta-Frau (2019)), specifically designed for analyzing 2D datasets. Building on the strengths of its one-dimensional counterpart, GradEn uniquely combines both symbolic representations and amplitude variations at each spatial point. This integration enhances GradEn’s discriminative power, allowing for more effective characterization of diverse image data and its classification applications. We begin by discussing the selection of parameters and the robustness of the GradEn method when applied to 2D white noise. Next, experimental validation of classification capability is carried out through systematic testing on simulated 2D datasets, which comprise colored noise and Mix processes. For logistic mapping, we employ the distance matrix from the recurrence plot method to convert the 1D time series into an image, enabling the identification of distinct images generated by different logistic parameters. Additionally, the computational cost is discussed to demonstrate the efficiency of GradEn. The method is also applied to real-world data: two real texture datasets are selected for classification tasks, and GradEn is used to detect and diagnose gear faults and railway corrugation in real-world signals after transforming these signals into matrix form. In both simulated and real-world experiments, the results of the proposed GradEn method are compared with other state-of-the-art 2D entropy algorithms, including $SampEn_{2d}$, $DispEn_{2d}$ and $PerEn_{2d}$. The classification results highlight the superiority and effectiveness of GradEn in distinguishing different types of textures and images, providing a novel perspective for analyzing the properties and structures of 2D data.
@@
The content of this paper is listed as follows: Section 2 provides a detailed explanation of the procedure for calculating GradEn. Section 3 presents the results of applying GradEn to simulated 2D data, demonstrating the effectiveness of the proposed method. In Section 4, we use real-world texture datasets and signals for classification using GradEn. Finally, Section 5 concludes the paper.
2 Gradient Entropy
@@
Slope entropy (Cuesta-Frau (2019)) is based on calculating the slope between consecutive points in a time series. This method offers a unique approach by incorporating both the symbolic pattern and the amplitude information within the series. When applied to 2D images, the concept of slope can be extended and replaced by the gradient of pixels in various directions.
@@
Consider an image $X=\{x_{i,j}\}_{i=1,2,...,H}^{j=1,2,...,W}$ with the size equals to $H\times W$. The sketch map of GradEn symbolization algorithm is shown in Fig. 1 and the procedure is listed as follows:
@@
3 Experiments in simuated data
3.1 Parameter Analysis
First of all, the selection of pixel block dimensions ($m*n$) in GradEn requires analysis, similar to the determination of embedding dimension $m$ in SloE. While the block size could theoretically be arbitrary, practical implementation must consider the structural characteristics of natural images. Smaller configurations such as $1*2$ or $2*1$ blocks yield only $55^{1}=$ 5 distinct gradient patterns(see Fig. 2(b)), which proves insufficient for effective information characterization. Conversely, larger configurations like $2*3$ or $3*2$ blocks generate $6255^{6}=15,$ 625 potential patterns (see Fig. 2(c)), creating dual challenges of computational complexity and noise sensitivity. Through evaluation, we identify the $2*2$ configuration as an optimal compromise: it produces $1255^{3}=$ 125 distinctive patterns - a quantity that maintains discriminative power while ensuring computational efficiency and noise robustness.
@@ 3.1 Parameter Analysis
Then, the thresholds $\delta$ and $\gamma$ should be selected carefully. Due to the z-score standardization, the range of $\delta$ is constrainted to $(\mu_{0.5},\mu_{0.75})$ and $\gamma$ falls within $(\mu_{0.75},\mu_{1})$, where $\mu_{\alpha}$ represents the $\alpha$ quantile of the normal distribution. Let $\delta=\mu_{a}$ and $\gamma=\mu_{b}$, a 100*100 2D whitenoise image with the size of 100*100 is then generated, and the GradEn values for $a$ changing from 0.51 to 0.74 and $b$ changing from 0.76 to 0.95 with 0.01 precision increments, are shown in Fig. 3. Given the high degree of irregularity in the white noise data, the entropy value is expected to be relatively large. Therefore, the parameter values in the yellow region ($a\in$ [0.53,0.6], $b\in$ [0.8,0.86]) in Fig. 3 are recommended. In this paper, $a=0.55$ and $b=0.8$ as the thresholds.
@@
3.2 Colored noise classification
@@ 3.3 Robust Analysis
3.4 Logistic map
The logistic map is a widely used chaotic model due to its ability to exhibit a variety of nonlinear behaviors as the parameter changesYang and Liao (2018). The formulation of the logistic map is given by:
@@ 3.4 Logistic map
where $a$ is the control parameter that determines the degree of periodicity or chaos in the sequence. To assess the ability of GradEn to distinguish between different states in the logistic map, we transform the logistic sequence into a distance matrix, similar to the preprocessing algorithm used in recurrence plots. Given the sequence $X_{i},i=1,2,...,N$, first perform phase space reconstruction with an embedding dimension $m$, resulting in a series of vectors $x(i)=(X_{i},X_{i+1},...,X_{i+m-1})$ for $i=1,2,...,N-m+1$. Next, construct the distance matrix $D=D_{i,j}$, where $i=1,2,...,N-m+1$ and $j=1,2,...,N-m+1$, with each element defined as $D_{i,j}=dist(x(i),x(j))$. The Euclidean distance is used in this algorithm. Once the matrix $D$ is constructed, it is then processed using the 2D entropy method to calculate the results.
@@ 3.5 Computation cost
Table 1: The Hedges’ g effect size results for the four methods across 15 EEG channels
@@
4 Experiments in real-world data
4.1 Texture Classification
Texture analysis is a key research area in image processing, focused on extracting the characteristics of various patterns within images. In this section, two texture datasets are used to evaluate the classification performance of the proposed $GradEn$ method. We begin by conducting experiments on the Kylberg texture dataset (Kylberg (2011)), which includes five selected texture categories: cushion, floor, scarf, screen, and stone. Each category contains 160 samples, and examples of the images are shown in Fig. 9. To standardize image size and ensure computational efficiency, all images are downsampled to 128×128. The boxplot results for $DistrEn_{2d}$ ($M=64$ and $M=128$), $PerEn_{2d}$, and $GradEn$ are presented in Fig. 10. As shown in Fig. 10(a) and (b), $DistrEn_{2d}$ fails to effectively distinguish between the five texture types, particularly for cushion, floor, and scarf. Similarly, $PerEn_{2d}$ ([Fig. 10(c)]) struggles to differentiate between cushion vs. scarf and floor vs. screen. In contrast, $GradEn$ demonstrates a clear ability to differentiate all five textures, as shown in [Fig. 10(d)], highlighting the superiority of the $GradEn$ method.
@@ 4.1 Texture Classification
Additionally, the Describable Texture Dataset (DTD) (Cimpoi et al. (2013)) is used for the classification task. Four texture types—knitted, polka, wrinkled, and zigzagged—are selected, with 120 samples for each type and the examples are shown in Fig. 11. All images are grayscaled processed by taking the means of RGB values and also downsampled to 128×128 for consistency. The boxplot results for the four entropy algorithms are presented in Fig. 12. It is evident that only $GradEn$ is able to distinguish between the four texture groups to some extent, while the other three methods fail to differentiate the texture types effectively. In conclusion, both two experiments showcase the effectiveness of $GradEn$ for detecting the characteristic of texture images.
@@
4.2 Gear fault diagnosis
Mechanical fault diagnosis plays a crucial role in industrial engineering, and numerous methods have been proposed by researchers in recent years (Xi et al. (2021); Zheng and Pan (2020); Shi et al. (2022); Sharma and Parey (2016)). In this section, gear fault data from the University of Connecticut are used for classification. The vibration signals are collected from a benchmark two-stage gearbox with replaceable gears. These signals are measured using an accelerometer and recorded through a dSPACE system, with a sampling frequency of 20 kHz. Four fault conditions—healthy, missing tooth, spalling, and chipping rip—are selected for the classification task (The examples are shown in Fig. 13). Each fault type consists of 104 samples. The first 150 points of each sample are used to construct the distance matrix with an embedding dimension of $m=3$. The boxplot results for classifying the four gear faults are shown in Fig. 14. As seen in the figure, $DistrEn_{2d}$ fails to effectively distinguish between the different fault types. $PerEn_{2d}$ struggles to differentiate between healthy vs. missing tooth and spalling vs. chipping rip. In contrast, only the $GradEn$ method is able to distinguish all four gear fault types based on entropy values, demonstrating its capability to characterize images generated by engineering signals.
@@ 4.2 Gear fault diagnosis
4.3 Railway Corrugation Diagnosis
The detection and diagnosis of railway corrugation is one of the most important technology part for keeping the security of train operation (Li and Shi (2020); Zhang and Shang (2022); Xu et al. (2023)). In this section, the acceleration signals collected from the sections where corrugation exists of heavy haul railway are applied to verify the detection ability of $GradEn$. The signals in this experiment are provided by School of Automation and Intelligence, Beijing Jiaotong University. Fig. 15(a) and Fig. 15(b) gives the two acceleration signals collected from the left and right axes of the railway section. The corrugation sections are marked by red frame in the picture. In our experiment, the signals with the length of $N=4000$ in both corrugation section (red frame) and normal sections (blue frame) are selected. The sliding window with $N=150$ with each time moves 10 points from the beginning to the end of the time series section. Thus, totally 385 samples of sections from both corrugation and normal signals are obtained. Then, all signals are transformed to distance matrix with embedding dimension $m=3$. The boxplot results of three entropy methods ($DistrEn_{2d}$, $PerEn_{2d}$, $GradEn$) for distinguishing corrugation and normal signals in both left and right axes are shown in Fig. 15(c-e) and (f-h). Moreover, the coresponding hedge effect size are calculated and given in each picture. The figures and numerical values manifest that $GradEn$ holds the best distinguishing results in both left and right axes signals, which proves that $GradEn$ method provides an effective tool for railway diagnosis research.
@@ 5 Conclusion
6 Statement
The paper is under consideration at Pattern Recognition Letters.
[…]